# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 01 Mar 2024 v4

@@ Appendix A Soft-thresholding SalUn
Appendix B Additional Experimental Details and Results
B.1 Additional training and unlearning settings
@@
For the Retrain method, training is conducted over $182$ epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at $0.1$. For both FT and RL, training spans 10 epochs with a fixed learning rate of $0.1$. GA’s training settings involve a 5-epoch learning rate search within the interval $10^{-6},10^{-4}]$. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range $[1,20]$. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range $10^{-6},10^{-4}]$, while maintaining a constant learning rate of 0.1. For the BS method, the FGSM step size is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval $10^{-6},10^{-4}]$. Lastly, for SalUn and SalUn-soft, we trained for 10 epochs, searching for learning rates in the range $10^{-4},10^{-2}]$ and sparsity ratios in the range $[0.2,0.6]$.
MU for image generation
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of $10^{-4}$. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of $10^{-4}$, $\alpha$ set to $10^{-3}$ and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
For SD, the unlearning settings are as follows: For SalUn, it undergoes training for 5 epochs with Adam using a learning rate of $10^{-5}$. The $\alpha$ value is set at 0.5, with a batch size of 8. The sparsity of weight saliency is again set at 50%. The sampling settings involve the use of DDIM, 100 time steps, and a conditional scale of 7.5.
@@
SalUn
@@
Additionally, Fig. A2 shows the examples of generated images when forgetting the class ‘airplane’. Opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A3, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a grid search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
B.2 Additional classification results
@@
FT + $\mathbf{m}_{\mathrm{S}}$ 4.13 (1.11) 98.35 (1.65) 92.66 (1.60) 8.40 (4.48) 2.21 2.38 7.92 (0.01) 96.49 (3.51) 88.50 (3.22) 12.87 (6.42) 3.29 1.34
RL 7.61 (2.37) 99.67 (0.33) 92.83 (1.43) 37.36 (24.48) 7.15 2.64 4.80 (3.11) 99.55 (0.45) 91.31 (0.41) 41.95 (22.66) 6.66 2.65
RL + $\mathbf{m}_{\mathrm{S}}$ 1.55 (3.69) 99.88 (0.12) 93.93 (0.33) 13.28 (0.41) 1.13 2.66 5.85 (2.06) 97.17 (2.83) 89.45 (2.27) 19.79 (0.50) 1.92 2.68
@@
B.3 Additional generation results
@@
Appendix C Broader Impacts and Limitations

### 04 Apr 2024 v5

@@ Appendix B Soft-thresholding SalUn
Appendix C Additional Experimental Details and Results
C.1 Additional training and unlearning settings
@@
For the Retrain method, training is conducted over $182$ epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at $0.1$. For both FT and RL, training spans 10 epochs within the interval $10^{-3},10^{-1}]$. GA’s training settings involve a 5-epoch learning rate search within the interval $10^{-5},10^{-3}]$. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range $[1,20]$. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range $10^{-6},10^{-4}]$, while searching for the learning rate within the range $10^{-3},10^{-1}]$. For the BS method, the step size of fast gradient sign method(FGSM) is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval $10^{-6},10^{-4}]$. Lastly, for SalUn and SalUn-soft, we trained for 10 epochs, searching for learning rates in the range $5*10^{-4},5*10^{-2}]$ and sparsity ratios in the range $[0.1,0.9]$.
@@
MU for image generation
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of $10^{-4}$. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of $10^{-4}$, $\beta$ set to $10^{-3}$ and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
For SD, the unlearning settings are as follows: For SalUn, it undergoes training for 5 epochs with Adam using a learning rate of $10^{-5}$. The $\beta$ value is set at 0.5, with a batch size of 8. The sparsity of weight saliency is again set at 50%. The sampling settings involve the use of DDIM, 100 time steps, and a conditional scale of 7.5.
Sparsity Ratio of Saliency Mask ($\mathbf{m}_{\mathrm{S}}$)
@@
Additionally, Fig. A2 shows the examples of generated images when forgetting the class ‘airplane’. Opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A3, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a grid search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a higher sparsity ratio is preferred. This is due to the fact that the larger forgetting data amount can exert the greater impact on the model, necessitating the higher saliency sparsity.
C.2 Additional classification results
@@
FT + $\mathbf{m}_{\mathrm{S}}$ 5.33 (0.09) 96.06 (3.94) 89.89 (4.37) 11.82 (1.06) 2.37 2.38 10.09 (2.18) 93.82 (6.18) 86.56 (5.16) 16.84 (2.45) 3.99 1.34
RL 7.61 (2.37) 99.67 (0.33) 92.83 (1.43) 37.36 (24.48) 7.15 2.64 7.61 (0.30) 99.67 (0.33) 92.83 (1.11) 37.36 (18.07) 4.95 2.65
RL + $\mathbf{m}_{\mathrm{S}}$ 2.85 (2.39) 99.62 (0.38) 93.93 (0.33) 14.39 (1.51) 1.15 2.66 7.75 (0.16) 94.28 (5.72) 89.29 (2.43) 16.99 (2.30) 2.65 2.68
@@
C.3 Additional generation results
@@
Appendix D Broader Impacts and Limitations